# Narrow line AGN selection in CEERS: spectroscopic selection, physical properties, X-ray and radio analysis

### v1

@@ 3.2 UV diagnostic
In this section, we will focus on the diagnostic diagrams involving rest-frame UV lines (Mingozzi et al., 2024; Mascia et al., 2024; Scholtz et al., 2023a; Feltre et al., 2016).
@@
In particular, we considered the UV diagnostic diagram C3He2-C43 and the diagnostic diagrams involving the so-called high-ionization emission lines, i.e. N iv] 1486, N v 1243, [Ne iv] 2425.
@@
The C3He2-C43 diagnostic diagram reported in Fig. 5 allows us to select one single NLAGN (CEERS-613), whose selection is based on a clear detection of the He ii 1640 line, that places the source well into the AGN-dominated region according to the demarcation lines presented in Hirschmann et al. (2022). The detection of the UV lines involved in this diagram is challenging with only min of JWST exposure (the average on-source exposure time of the CEERS survey). Also in Scholtz et al. (2023a) the number of detections of these lines is low, even if the exposure time per target of the JADES survey was times longer than the CEERS observations.
@@
The high-ionization emission lines are characterized by high-photoionization energies that can be produced more probably in AGN, due to their harder ionizing radiation, and therefore can be used as indicators of the presence of an AGN, even if hidden. In the diagnostics diagrams involving these lines, we base the identification of NLAGN on the position of the sources compared to the distribution of AGN and SFG photoionization models of Gutkin et al. (2016) and Feltre et al. (2016). In the top panel of Fig. 6 we show the diagnostic diagram involving the N iv] 1486 emission line, which has an ionization potential of 47eV. This diagnostic allows us to select two NLAGN (CEERS-496, CEERS-1019), based on their extreme position with respect to the distribution of both SFG and AGN photoionization models. However, given their unclear detection of the He ii 1640 line in both these sources, and given the uncertainties in the extension of the SFG region in the diagnostic, we decided to mark these NLAGNs only as candidates. In particular, the source in the upper part of the diagram is CEERS-1019, whose AGN nature has already been discussed in multiple works in the literature (Mascia et al., 2024; Larson et al., 2023), some supporting the presence of an AGN and some discarding it (Tang et al., 2023). In the middle panel of the same figure is represented the diagnostic diagram involving the N v 1240 emission line (ionization potential of 77eV). In this case, we could select only one source with a N v 1240 detection, CEERS-23, that again we marked as a candidate due to the undetection of the He ii 1640 line. In the last panel of Fig. 6, we show the [Ne iv] 2424 emission line diagnostic, whose ionization potential is 63eV. We detect the [Ne iv] 2424 line in three different sources (CEERS-1029, CEERS-13061, CEERS-4210), and all of them occupy a region of the diagram covered only by the AGN models, even considering the upper limits in He ii 1640. In this diagnostic diagram, where the [Ne iv] and [Ne iii ]lines are apart one from the other, the dust correction would move the points towards the right, to even more AGN-extreme values.
@@
Furthermore, we detect the [Ne v] 3426 emission line (97eV of ionization potential) in one source (CEERS-8299, at ) that we marked as AGN given the extremely high-ionization energy required to produce this line. Even if at high- it has been proposed that the [Ne v] 3426 could also be associated with SF processes Cleri et al. (2023), it was always related to AGN activity at Mignoli et al. (2013); Barchiesi et al. (2024).
Overall we selected 52 NLAGN among the initial 217 sources.
@@ 3.3 Stack of the AGN and non-AGN samples
In each of the diagnostic diagrams reported in Sec. 3, we also report the line ratios of the stacked spectra of all the selected NLAGN and of the non-AGN selected sources, derived as presented in Sec. 2.3 and shown in Fig. 1. The two stacked spectra were then fitted to measure emission line fluxes using Qubspec, following the same procedure as for all the other single spectra and described in Sec. 2.2.
As it is possible to see from the BPT (R3N2) diagram, the position of the stacked NLAGN and the non-AGN sources are very close together, with AGN having a slightly larger R3 line ratio. This clearly shows how the BPT diagram and its traditional demarcation lines are no longer useful in effectively separating the AGN population from SFG, probably even for relatively low redshifts (). We will further discuss this point in Sec. 4.1. On the contrary, in the R3S2 diagnostic diagram (VO87), the line ratios of the stacked NLAGN spectrum clearly lie in the AGN region of the diagnostic, while the non-AGN one is in the SFG-dominated region. The R3S2 line ratio seems to be still informative for the AGN selection also at (relatively) high-, as we will discuss in Sec. 4.1. The same applies to the R3O1 diagnostic diagram. As for the He2N2 diagram, we note that while the NLAGN stacked spectrum shows a clear detection of the He ii 4686 line, which places the final NLAGN stack clearly in the AGN region of the diagnostic, for the non-AGN sample the He ii 4686 line is not detected and the line ratios fall in the SFG region, as expected. This is also a strong point in favor of the goodness of our NLAGN selection since the He ii 4686 line was detected with high-enough S/N only in 5/52 of the NLAGN, 4 or them at . The fact that this line is clearly detected in the NLAGN spectral stack and that the stack line ratios fall in the AGN region of the diagnostic means that signatures of AGN emission are widely present among our NLAGN sample.
In both the diagnostic diagrams involving the [O iii] 4363 auroral line, NLAGN are characterized by a larger ( dex) [O iii] 4363/H line ratio compared to the non-AGN, while the two have a similar O3O2 or Ne3O2 line ratios, meaning that the NLAGN selection is mainly driven by a stronger [O iii] 4363 auroral line in AGN, as discussed in Mazzolari et al. (2024).
We did not put the results of the stack in the rest-UV diagnostic diagrams. This is because in the case of the C3He2-C43 diagnostics, the He ii 1640 and C iv lines are available only for 33/155 non-AGN and 7/52 NLAGN, and both these lines are not clearly detected in none of the two stacked spectra. This is expected given the faintness of these lines and the observing time of the CEERS survey, and indeed this diagnostic was able to identify only 1 NLAGN. The same goes for the high-ionization emission line diagnostics: none of the high-ionization emission lines involved in the diagnostics in Fig. 6 is clearly detected in our stack.
@@ 4 Discussion
The IDs and the main physical properties (that will be presented and discussed in the next Sections) of all the selected 52 NLAGN are reported in Table 3.
@@ 4.1 Comparison of AGN selection methods
In this section, we discuss the effectiveness of the different diagnostic diagrams reported in Sec. 3. For the discussion below, the term effectiveness refers to the ability of a diagnostic diagram to efficiently select (i.e. isolate in the diagram) NLAGN among all the considered sources.
@@
As presented in the top panel of Fig. 7, most of the NLAGN are selected based on only one single diagnostic diagram (35/52, the 65%), 13 are selected in two different AGN diagnostics, while 4 are selected based on more than two diagnostics diagrams. The fact that most of the NLAGN are selected based on one single NLAGN diagnostic is not surprising. The NLAGN diagnostic diagrams we used for our selection involve emission lines that cover a wide wavelength range (from rest-UV to rest-optical), over which the spectral features of sources at can vary a lot depending on many parameters (metallicities, stellar population in the host galaxy, radio or X-ray emission, etc.). Given this, different NLAGN diagnostic diagrams are intrinsically sensitive to different AGN properties and, in turn, can become ineffective (and therefore less useful in separating AGN from SFG) under different conditions.
@@
Looking at the lower panel of Fig. 7, we see the distribution of the sources selected by the different diagnostic diagrams with respect to their redshift and their selection diagram. The diagnostic diagram that selected the larger number of the AGN is the R3S2, in particular at , but with a non-negligible number of NLAGN also selected at . The effectiveness of the R3S2 (but also of the R3O1) diagnostic diagram in the low-metallicity (and therefore high-) regime was already pointed out in some recent works investigating the effectiveness of traditional NLAGN diagnostic diagrams to select AGN among low-metallicity dwarf galaxies (Polimera et al., 2022). In particular, they show that the S2 and the O1 line ratio are less metallicity sensitive and more successful in identifying AGN in these dwarf environments (more similar to those at high-) than the R3N2 diagnostic. The lower effectiveness of the R3N2 at low metallicities (hence at high redshift) is an obvious consequence of the fact that the nitrogen abundance scales about quadratically with metallicity (at ), hence the [NII]/H ratio drops strongly, even if irradiated by the energetic radiation of an AGN. We also notice that [N ii] and [S ii] get redshifted out from the reddest grating of the spectra at , while the [O i] gets redshifted out at .
@@
In general, at the most effective AGN selection techniques are the R3S2, N2He2, and R3N2, while the large majority of the AGN at are selected based on the auroral line diagnostics (both M1 and M2 ), or based on the detection of high-ionization emission lines. This supports the picture in which the traditional AGN diagnostic diagrams (in particular the R3N2) become less effective at high-, mainly due to metallicity-related effects, while the detection of high-ionization emission lines and the new [O iii] 4363-based AGN diagnostics prove to be effective in selecting AGN also in the early Universe, reinforcing the results pointed out by Maiolino et al. (2023), Übler et al. (2023) and Scholtz et al. (2023a).
@@ 4.2 Selection of X-ray and radio sources in the diagnostics
In this section, we want to summarize the results related to the X-rays and radio sources involved in our analysis. Among the seven X-ray sources with a spectrum in the CEERS MR sample, two sources (CEERS-3050 and CEERS-3061) are at , and therefore, the rest-frame optical and UV lines are not available in JWST spectra. These sources were classified as SFG by the X-ray spectral analysis performed in Buchner et al. (2015). Among the other five X-ray sources, four are selected as NLAGN in at least one of the diagnostic. In particular, the one at highest redshift (CEERS-2808, ) is selected as NLAGN only in the R3S2 diagnostic diagram and not in the R3N2 (considering the more conservative demarcation line reported in Scholtz et al., 2023a). The only three X-ray sources selected as NLAGN in the BPT diagram are CEERS-2919, CEERS-2904, and CEERS-2900, all at , the first two also showing a broad H emission line (see Appendix A). The only X-ray source not selected in any of the diagnostics is CEERS-2989, at , whose spectrum shows only clear H and [N ii] detections, and therefore, we could not define the R3 line ratio to place it in the above-mentioned diagnostics. It is worth noting that the lack of [O iii] 5007 and H detections in this source is probably due to high obscuration levels. Considering the analysis in Sec. 4.6, we found a lower limit to the value of mag. The large AGN obscuration of this source is further supported by the Compton-thick obscuration level derived by the X-ray analysis performed in Buchner et al. (2015). However, this source shows a faint broad H component, as reported in Fig. 17.
The two radio-detected sources are CEERS-2900 (that is also X-ray detected) and CEERS-3129, both selected in the R3N2 diagram as well as in the R3S2 one. They also both show strong [N ii] emission, in particular CEERS-3129, probably indicative of a shock (Allen et al., 2008; Nesvadba et al., 2017). Furthermore, CEERS-3129 shows a significant broad H component, but no indication of X-ray emission, as we will further explore in Sec. 4.9.
@@ 4.3 Comparison with previous CEERS AGN selections
Calabrò et al. (2023) attempted to select AGN among the CEERS sources with a MR spectrum using near-infrared emission line diagnostics. To do so, they restricted the redshift range of the sources at , therefore considering only 65 single sources. Using the R3N2 and R3S2 diagnostic diagrams, the authors selected 8 NLAGN that were classified as NLAGN also based on at least one NIR diagnostic, i.e. CEERS-2919, CEERS-3129, CEERS-2904, CEERS-2754, CEERS-5106, CEERS-12286, CEERS-16406, and CEERS-17496. The first three were marked as BLAGN, as we also found in Appendix 17. Five of the sources selected as NLAGN by Calabrò et al. (2023) are also selected in our work, while the other three sources were not selected because of the more conservative demarcation lines used in the R3S2 and R3N2 diagnostics or because the [S ii] line was not considered detected with enough significance by our fit. The authors also selected 5 NLAGN using near-infrared diagnostic diagrams that were classified as SFG based on the optical diagnostic diagrams, i.e. CEERS-2900, CEERS-8515, CEERS-8588, CEERS-8710, CEERS-9413. Among these, the only one that is classified as NLAGN also in our selection is CEERS-2900; the other three were selected as SFG also by our optical diagnostics (even if CEERS-8588 shows a tentative [Ne v] line emission, but we conservatively decided to mark it as a non-detection).
In Davis et al. (2023) the authors investigate the presence of extreme emission-line galaxies (EELG) at using JWST NIRCam photometry in the CEERS program. They used a method to photometrically identify EELGs with H + [O iii] or H emission of observed-frame equivalent width . Among the photometrically selected EELGs there are 39 sources with a NIRSpec PRISM or MR spectrum, 25 with a match in our MR parent sample. Among these, 6 are selected as NLAGN in our work, i.e., , supporting the non-negligible AGN contamination in the photometric selection of this kind of sources, as already shown also by other works (Amorín et al., 2015).

### v3

@@ 3.2 UV diagnostic
In this section, we focus on the diagnostic diagrams involving rest-frame UV lines (Mingozzi24; Mascia24; Scholtz23b; Feltre16).
@@
In particular, we considered the UV diagnostic diagram C3He2-C43 and the diagnostic diagrams involving the so-called high-ionization emission lines, N iv] 1486, N v 1243, and [Ne iv] 2425.
@@
The C3He2-C43 diagnostic diagram reported in Fig. 5 allows us to select one single NLAGN (CEERS-613), whose selection is based on a clear detection of the He ii 1640 line, which places the source well into the AGN-dominated region according to the demarcation lines presented in Hirschmann22. The detection of the UV lines involved in this diagram is challenging with only min of JWST exposure (the average on-source exposure time of the CEERS survey). Also in Scholtz23b the number of detections of these lines is low, even if the exposure time per target of the JADES survey was times longer than the CEERS observations.
@@
The high-ionization emission lines are characterized by high-photoionization energies that are more likely to be produced in AGNs, due to their harder ionizing radiation, and therefore can be used as indicators of the presence of an AGN, even if hidden. In the diagnostics diagrams involving these lines, we base the identification of NLAGN on the position of the sources compared to the distribution of AGN and SFG photoionization models of Gutkin16 and Feltre16. In the top panel of Fig. 6 we show the diagnostic diagram involving the N iv] 1486 emission line, which has an ionization potential of 47eV. This diagnostic allows us to select two NLAGN (CEERS-496, CEERS-1019), based on their extreme position with respect to the distribution of both SFG and AGN photoionization models. However, given the unclear detection of the He ii 1640 line in both these sources, and given the uncertainties in the extension of the SFG region in the diagnostic, we decided to mark these NLAGNs only as candidates. In particular, the source in the upper part of the diagram is CEERS-1019, whose AGN nature has already been discussed in multiple works in the literature, some supporting the presence of an AGN based on a low-significance broad H detection (Mascia24; Larson23) and some discarding it (Tang23) and attributing its spectral properties to exotic SF. In the middle panel of the same figure is represented the diagnostic diagram involving the N v 1240 emission line (ionization potential of 77eV). In this case, we could select only one source with a N v 1240 detection, CEERS-23, which we again marked as a candidate due to the undetection of the He ii 1640 line. In the last panel of Fig. 6, we show the [Ne iv] 2424 emission line diagnostic, whose ionization potential is 63eV. We detect the [Ne iv] 2424 line in three different sources (CEERS-1029, CEERS-13061, CEERS-4210), and all of them occupy a region of the diagram covered only by the AGN models, even considering the upper limits in He ii 1640. In this diagnostic diagram (where the [Ne iv] and [Ne iii ]lines are apart one from the other) the dust correction would move the points toward the right, to even more AGN-extreme values.
@@
Furthermore, we detect the [Ne v] 3426 emission line (97eV of ionization potential) in one source (CEERS-8299, at ) that we marked as AGN given the extremely high-ionization energy required to produce this line. Even if at high- it has been proposed that the [Ne v] 3426 could also be associated with SF processes (Cleri23), it was always related to AGN activity at Mignoli13; Barchiesi24. Overall, we selected 52 NLAGNs among the initial 217 sources.
@@ 3.3 Stack of the AGN and non-AGN samples
In each of the diagnostic diagrams reported in Sec. 3, we also report the line ratios of the stacked spectra of all the selected NLAGN and of the non-AGN selected sources, derived as presented in Sec. 2.3 and shown in Fig. 1. The two stacked spectra were then fit to measure emission line fluxes using Qubspec, following the same procedure as for all the other single spectra and described in Sec. 2.2.
As it is possible to see from the R3N2 diagram, the position of the stacked NLAGN and the non-AGN sources are very close together, with AGN having a slightly larger R3 line ratio. This clearly shows how the BPT diagram and its traditional demarcation lines are no longer useful in effectively separating the AGN population from SFG, probably even for relatively low redshifts (). We further discuss this point in Sec. 4.1. On the contrary, in the R3S2 diagnostic diagram, the line ratios of the stacked NLAGN spectrum clearly lie in the AGN region of the diagnostic, while the non-AGN one is in the SFG-dominated region. The R3S2 line ratios seem to be still informative for the AGN selection also at (relatively) high-, as we discuss in Sec. 4.1. The same applies to the R3O1 diagnostic diagram. As for the He2N2 diagram, we note that while the NLAGN stacked spectrum shows a clear detection of the He ii 4686 line, which places the final NLAGN stack clearly in the AGN region of the diagnostic, for the non-AGN sample the He ii 4686 line is not detected and the line ratios fall in the SFG region, as expected. This is also a strong point in favor of the goodness of our NLAGN selection since the He ii 4686 line was detected with high-enough S/N only in 5/52 of the NLAGN, 4 of them at . The fact that this line is clearly detected in the NLAGN spectral stack and that the stack line ratios fall in the AGN region of the diagnostic means that signatures of AGN emission are widely present among our NLAGN sample.
In both the diagnostic diagrams involving the [O iii] 4363 auroral line, NLAGN are characterized by a larger ( dex) [O iii] 4363/H line ratio compared to the non-AGN, while the two have a similar O3O2 or Ne3O2 line ratios, meaning that the NLAGN selection is mainly driven by a stronger [O iii] 4363 auroral line in AGN, as discussed in Mazzolari24b.
We did not put the results of the stack in the rest-UV diagnostic diagrams. This is because the He ii 1640 and C iv lines are available only for 33/155 non-AGN and 7/52 NLAGN, and both these lines are not significantly detected in none of the two stacked spectra. This is expected given the faintness of these lines, the possible stronger effect of obscuration at rest-UV wavelengths, and the observing time of the CEERS survey. The same goes for the high-ionization emission line diagnostics: none of the high-ionization emission lines involved in the diagnostics in Fig. 6 is clearly detected in our stack.
@@ 4 Discussion
The IDs and the main physical properties (that will be presented and discussed in the next Sections) of all the 52 NLAGN selected in Sec. 3 are reported in Table 8.
@@ 4.1 Comparison of AGN selection methods
In this section, we discuss the effectiveness of the different diagnostic diagrams reported in Sec. 3. For the discussion below, the term effectiveness refers to the ability of a diagnostic diagram to efficiently select (i.e., isolate in the diagram) NLAGN among all the considered sources.
@@
As presented in the top panel of Fig. 7, most of the NLAGN are selected based on only one single diagnostic diagram (35/52, the 65%), 13 are selected in two different AGN diagnostics, while 4 are selected based on more than two diagnostics diagrams. The approach chosen for the NLAGN selection in this work clearly favors the high-completeness instead of the high-purity criterium, selecting as NLAGN all the sources showing the dominance of the AGN photoionization in at least one diagnostic diagram. However, the fact that most of the NLAGN are selected based on one single NLAGN diagnostic is not surprising. The NLAGN diagnostic diagrams used in our work involve emission lines that cover a wide wavelength range (from rest-UV to rest-optical), over which the spectral features of sources at can vary a lot depending on many parameters (metallicities, stellar population in the host galaxy, radio or X-ray emission, etc.). Given this, different NLAGN diagnostic diagrams are intrinsically sensitive to different AGN properties and, in turn, can become ineffective (and therefore less useful in separating AGNs from SFGs) under different conditions. Additionally, given the wide redshift range of the sources analyzed in this work, not all of them have in their observed spectra all the lines used for the diagnostic reported in Sec.3.1,3.2, mostly because lines might be redshift out from the observed wavelength range or fall into a spectral gap.
@@
Looking at the lower panel of Fig. 7, we see the distribution of the sources selected by the different diagnostic diagrams with respect to their redshift. The diagnostic diagram that selected the larger number of the AGN is the R3S2, in particular at , but with a non-negligible number of NLAGN also selected at . The effectiveness of the R3S2 (but also of the R3O1) diagnostic diagram in the low-metallicity (more frequent at high-) regime was already pointed out in some recent works investigating the effectiveness of traditional NLAGN diagnostic diagrams to select AGN among low-metallicity dwarf galaxies (Polimera22). In particular, they show that the S2 and the O1 line ratios are less metallicity-sensitive and more successful in identifying AGN in these dwarf environments (more similar to those at high-) than the R3N2 diagnostic. The lower effectiveness of the R3N2 at low metallicities (hence at high redshift) can be a consequence of the nitrogen production channel (Henry00) that can determine nitrogen abundance to scale about quadratically with metallicity (at ), hence the [NII]/H ratio drops strongly. On the contrary, Sulfur and oxygen aboundances, being directly produced by massive stars through -processes, can be less dependent on metallicity (Dopita13). We also notice that [N ii] and [S ii] get redshifted out from the reddest grating of the spectra at , while the [O i] gets redshifted out at .
We further note that [O i] emission can sometimes be excited also by shocks, producing high [O i] /H ratios not necessarily driven by a dominant AGN ionization. However, shocks in SFG are generally sub-dominant, as evidenced by the distribution of SDSS galaxies in the BPT and by works selecting this kind of population among SDSS SFG (e.g. Alatalo16, found that shock-dominated SFG are of the SDSS DR7 SFG). For this reason, it would be unlikely that such a large fraction of SFG at higher redshift would exhibit shock-driven emission.
Shocks can also be an important source of ionizing power in mergers (Medling15) and, of course, in AGN (Best00). While, after a visual inspection, we did not find strong evidence of ongoing mergers in the majority of the sources selected as NLAGN by the R3O1 diagnostic, we note that 8 out of the 12 NLAGN selected using the R3O1 line ratios are also selected as NLAGN in other diagnostic diagrams, supporting the scenario in which for most of the sources the [O i] line should be driven by AGN photoionization.
@@
In general, at the most effective AGN diagnostic diagrams are the R3S2, N2He2, and R3N2, while the large majority of the AGN at are selected based on the auroral line diagnostics (both M1 and M2 ), or based on the detection of high-ionization emission lines. This supports the fact that traditional AGN diagnostic diagrams (in particular the R3N2) become less effective at high-, mainly due to metallicity-related effects, while diagnostic based on high-ionization emission lines or on the [O iii] 4363 line proved to be effective in selecting AGN also in the early Universe, as already pointed out by Maiolino23c, ubler23 and Scholtz23b.
@@ 4.2 Selection of X-ray and radio sources in the diagnostics
In this section, we want to summarize the results related to the X-rays and radio sources involved in our analysis. Among the seven X-ray sources with a spectrum in the CEERS MR sample, two sources (CEERS-3050 and CEERS-3061) are at , and their rest-frame optical and UV lines are not available in JWST spectra. These sources were classified as SFG by the X-ray spectral analysis performed in buchner15. Among the other five X-ray sources, four are selected as NLAGN in at least one of the diagnostics. In particular, the one at highest redshift (CEERS-2808, ) is selected as NLAGN only in the R3S2 diagnostic diagram and not in the R3N2 (considering the more conservative demarcation line reported in Scholtz23b). The only three X-ray sources selected as NLAGN in the BPT diagram are CEERS-2919, CEERS-2904, and CEERS-2900, all at , the first two also showing a broad H emission line (see Appendix A). The only X-ray source not selected in any of the diagnostics is CEERS-2989, at , whose spectrum shows only H and [N ii] detections and we could not define the R3 line ratio to place it in the above-mentioned diagnostics. It is worth noting that the lack of [O iii] 5007 and H detections in this source is probably due to high obscuration levels. Considering the analysis in Sec. 4.6, we found a lower limit to the ISM obscuration corresponding to mag. The large AGN obscuration of this source is further supported by the CTK obscuration level derived by the X-ray analysis performed in buchner15. However, this source shows a faint broad H component, as reported in Fig. 17.
The two radio-detected sources are CEERS-2900 (that is also X-ray detected) and CEERS-3129, both selected in the R3N2 diagram as well as in the R3S2 one. They also both show strong [N ii] emission, in particular CEERS-3129, which could be indicative of a shock (Allen08; Nesvadba17). Furthermore, CEERS-3129 shows a significant broad H component, but no indication of X-ray emission, as we further explore in Sec. 4.9.
@@ 4.3 Comparison with previous CEERS AGN selections
Calabro23 attempted to select AGN among the CEERS sources with a MR spectrum using near-infrared emission line diagnostics. To do so, they restricted the redshift range of the sources to , therefore considering only 65 sources. Using the R3N2 and R3S2 diagnostic diagrams, the authors selected 8 NLAGN that were classified as NLAGNs also based on at least one NIR diagnostic: CEERS-2919, CEERS-3129, CEERS-2904, CEERS-2754, CEERS-5106, CEERS-12286, CEERS-16406, and CEERS-17496. The first three were marked as BLAGN, as we also found in Appendix 17. Five of the sources selected as NLAGN by Calabro23 are also selected in our work, while the other three sources were not selected because of the more conservative demarcation lines used in the R3S2 and R3N2 diagnostics or because the [S ii] line was not considered detected with enough significance by our fit. The authors also selected five NLAGNs using near-infrared diagnostic diagrams that were classified as SFG based on the optical diagnostic diagrams: CEERS-2900, CEERS-8515, CEERS-8588, CEERS-8710, and CEERS-9413. Among these, the only one that is classified as NLAGN also in our selection is CEERS-2900; the other three were selected as SFG also by our optical diagnostics (even if CEERS-8588 shows a tentative [Ne v] line emission, but we conservatively decided to mark it as a non-detection).
In Davis23 the authors investigate the presence of extreme emission-line galaxies (EELG) at using JWST NIRCam photometry in the CEERS program. They used a method to photometrically identify EELGs with H + [O iii] or H emission of observed-frame equivalent width . Among the photometrically selected EELGs there are 39 sources with a NIRSpec PRISM or MR spectrum, 25 with a match in our MR parent sample. Among these, 6 are selected as NLAGNs in our work, in other words , supporting the non-negligible AGN contamination in the photometric selection of this kind of source, as has already been shown by other works (amorin15).